predict productso 0 NaBH4,420H فلم OCHZ 0 - H excess NH2NH2/H+ KOH/H2O/Heat ہ H
Answer #1

1.

Answer:

Sodium borohydride (NaBH4) is a mild reducing reagent that transforms aldehydes and ketones to the corresponding alcohol, pri

2.

Answer:

0 Н CH3 excess NH NH2, H KOH/H2O/Heat CH3 o H Product

Mechanism:

He OH HA NEMA H -p+H2M-NH2 -2H₂O LO+H₂N-NH₂ H N-NH 했 H Н H-OH Loon Huang Н I MIH H Hie N=N-H -2N2 EN-H H N=N-14 H OH H ento 6